Solve the differential equation
(tanxtany − e^ycosx) + y'sec^2 y = 0.
1
Expert's answer
2020-05-13T18:24:26-0400

the question is wrong.


the question should be:

tanx tany - ex cosx + y' sec2y = 0 equation1


solution: this equation is of the form:

f'(y) y' + P f(y) = Q where P and Q are functions of x.

here, P= tanx and Q = excosx , f(y)=tany


let tany=v such that sec2y. y' = v'

equation1 becomes v'+ vtanx = excosx equation2

this equation is linear in v and x

so, integrating factor= "\left(e^{\smash{\int P dx}}\right)" = "\left(e^{\smash{\int tanx dx}}\right)" = secx


equation2 becomes v.secx = "\int" ex cosx secs dx + constant'c'

v.secx = ex + c

tany.secx = ex +c